# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@ 3.3. Properties of classes of nearby representations
    The generic representation has rank
    $r_{gen}=\sum_{i}\left(\sum_{j}n^{+}_{ij}r^{+}_{ij}+\sum_{j}n^{-}_{ij}r^{-}_{ij}\right).$
@@
    The $i$-th nongeneric representation has rank
    $r_{i}=\sum_{j}n^{+}_{ij}R^{+}_{ij}+\sum_{l\neq i}\left(\sum_{j}n^{+}_{lj}r^{+}_{lj}+\sum_{j}n^{-}_{lj}r^{-}_{lj}\right).$
@@
We can also express the number of singularities at finite distance of the nearby representations:
@@
Let $(E,\nabla)$ of generic form, and $(E^{\prime},\nabla^{\prime})\in\operatorname{SL}_{2}(\mathbb{C})\cdot(E,\nabla)$. If $(E^{\prime},\nabla^{\prime})$ is in the $i$-th nongeneric class of representations, its number of singularities at finite distance is the number of distinct coefficients $b_{ij}$ among the circles in $N_{i}^{-}$.
@@
This is a consequence of item (2) of Prop. 2.14: the position of the singularity above which $F\cdot\langle q^{-}_{i,j}\rangle$ lies is determined by the coefficient $b_{ij}$. ∎
In particular, if $N_{i}^{-}=\emptyset$, then the $i$-th nongeneric class of representations only has a singularity at infinity. In general, the class of nearby representations all have different ranks and number of singularities.
Finally, we can also read from the generic tree when two nongeneric representations have the same forest.
@@
Keeping previous notations, for $i,j=1,\dots k$, with $i\neq j$, the fission forests $\mathbf{F}_{i}$ and $\mathbf{F}_{j}$ are isomorphic if and only if the $i$-th and $j$-th principal subtrees of the fission tree $\mathbf{T}$ are isomorphic.
@@
Since isomorphism classes of fission trees characterise admissible deformations, it is equivalent to show that the $i$-th and the $j$-th nongeneric forests are isomorphic if and only if $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are admissible deformations of each other. Let us first assume that $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are admissible deformations of each other. We have to be careful that in general if two irregular class $\Theta$ and $\Theta^{\prime}$ at infinity are admissible deformations of each other, it is not always true that so are their Fourier transform, because if $\langle q\rangle$ is a Stokes circle for $\Theta$ and $\langle q^{\prime}\rangle$ the Stokes circle for $\Theta^{\prime}$ corresponding to $\langle q\rangle$ under some admissible deformation, their slope may differ which may imply that $F\cdot\langle q\rangle$ and $F\cdot\langle q^{\prime}\rangle$ are not admissible deformations of each other. However here because all Stokes circles of $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are of slope $<2$ this cannot happen and we have that $F\cdot\Theta_{i}^{<2}$ and $F\cdot\Theta_{j}^{<2}$ are admissible deformations of each other. In turn, the irregular classes $\Theta_{i}^{<2}+\sum_{l\neq i}\Theta_{l}$ and $\Theta_{j}^{<2}+\sum_{l\neq j}\Theta_{l}$ are admissible deformations of each other, so the $i$-th and $j$-th representations coincide. Conversely, if the $i$-th and $j$-th classes of representations coincide, then in particular $\sum_{l\neq i}\Theta_{l}$ and $\sum_{l\neq j}\Theta_{l}$ are admissible deformations of each other, which implies that $\Theta_{i}$ and $\Theta_{j}$ are admissible deformations of each other, and the conclusion follows. ∎
Corollary 3.17.
Let $\bm{\breve{\Theta}}$ be an effective modified global irregular class. Then the number of isomorphism classes of fission forests among the forests of elements of the orbit $\operatorname{SL}_{2}(\mathbb{C})\cdot\bm{\breve{\Theta}}$ is equal to 1 plus the number of distinct isomorphism classes of fission trees among the principal subtrees of its generic fission tree.
@@
The classes of nearby representations arise from the action of $\operatorname{SL}_{2}(\mathbb{C})$ on connections on the affine line. However, as already mentioned in the introduction, there is a larger class of natural operations on connections: indeed one can consider Möbius transformations, which have the effect of changing the point at infinity in the identification $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$, as well as the operation consisting in tensoring a connection with an arbitrary rank one local system on a Zariski open subset of $\mathbb{P}^{1}$. This larger class of operations on connections on $\mathbb{P}^{1}$ plays an essential role in the extension due to Deligne-Arinkin [17, 3] of the Katz algorithm for rigid local systems to the case of irregular connections, which provides a way to reduce any irreducible rigid connection on $\mathbb{P}^{1}$ to the trivial rank one connection by repeated application of one of these operations. Since the symplectic $\operatorname{SL}_{2}(\mathbb{C})$ transformations don’t commute with these other operations in general, the latter do not preserve the diagram or the classes of nearby representations.
4. Basic representations for Painlevé moduli spaces
@@
In this section, as a set of examples illustrating our general discussion, we describe the basic representations of all Painlevé moduli spaces, obtained from the standard rank 2 Lax representation of each Painlevé equation. Recall that, while there were not first discovered in this way historically, all Painlevé equations can be obtained from isomonodromic deformations. Each Painlevé equation arises from isomonodromic deformations of connections with a specific type of singularity data, i.e. different Painlevé equations correspond to different nonabelian Hodge spaces. In the standard way to derive the Painlevé equations from isomonodromy [33], all Lax representations are of rank 2, and in all cases except for Painlevé VI, these Lax representations have at least one irregular singularity.
@@
For each case, starting from this standard Lax representation (in its abstract form, i.e. consisting in the data of a wild Riemann surface), we will draw the corresponding generic fission tree of minimal rank, describe the classes of nearby representations, and discuss when some nongeneric representations have the same forest. While the cases of Painlevé VI, V, IV all fall in the simply-laced case and were discussed by Boalch in [7], it is still worth discussing them briefly from the point of view of generic fission trees. On the other hand, describing all basic representations for Painlevé III, II, I as well as for the moduli spaces corresponding to degenerate versions of Painlevé III really requires using our general framework. Doing this, we recover among the classes of nearby representations many known alternative Lax representations for various Painlevé equations, e.g. from [25, 34, 35] and the other representations can be viewed as new abstract Lax representations.
@@
Notice that there is the following subtlety: since taking the Fourier transform doesn’t always commute with doing admissible deformations, different connections having the same fission forest in general do not have the same classes of nearby representations, nor the same diagram. In the cases where this happens, among all possible wild Riemann surfaces with boundary data having the fission forest of the standard Painlevé Lax representation, we choose one such that its generic representation has minimal rank. Making this choice is necessary to allow us to recover several known alternative Lax representations for Painlevé equations and to recover the diagrams for Painlevé equations obtained in [14], all equal or related to affine Dynkin diagrams. Other choices would lead to different diagrams, and for each choice all nearby representations could also be explicitly determined, yielding even more abstract Lax representations, typically of higher rank.
To draw the fission forests and the generic fission tree, we follow the conventions of [11]. The root of the trees is indicated by a black square vertex and the heights of vertices of the trees are indicated to the left of the figure. In [11] it was necessary to introduce different types of vertices, represented by different colors (black or white): we haven’t recalled the corresponding definitions here since this distinction doesn’t play a role for our present purposes, but will still represent the vertices with their right colours for the sake of consistency. The multiplicity of the leaves (i.e. the rank the corresponding conjugacy classes) is equal to 1 unless indicated otherwise.
@@
To draw the diagrams, we follow the same conventions as in [18]: unless otherwise specified the multiplicity of each vertex is 1, negative edges or loops are drawn in dashed lines, and for multiple edges/loops the multiplicity in indicated at their middle.
@@
The generic fission trees as well as the rank and number of singularities of each class of basic representations for each Painlevé moduli space are listed below. In the next paragraphs, we explain in more detail how to obtain these classes of representations starting from the standard Lax representations.
@@
22225/3535/35 / $34/$ 311112/3232/32 / $31/$ 3
Rank Number of singularities
Generic rep. 3 1
Nongeneric rep. 2 1
@@
111122223/2323/23 / 2
Rank Number of singularities
Generic rep. 3 1
1st nongeneric rep. 2 1
2nd nongeneric rep. 2 2
@@
111122221/2121/21 / 2
Rank Number of singularities
Generic rep. 4 1
1st nongeneric rep. 2 3
2st nongeneric rep. 2 2
@@
111122221/2121/21 / 2
Rank Number of singularities
Generic rep. 4 1
Nongeneric reps. 2 2
@@
111122221/2121/21 / 21/3131/31 / 3
Rank Number of singularities
Generic rep. 5 1
1st nongeneric rep. 2 2
2st nongeneric rep. 3 2
@@
22221111
Rank Number of singularities
Generic rep. 3 1
Nongeneric reps. 2 2
@@
11112222
Rank Number of singularities
Generic rep. 4 1
Nongeneric reps. 2 3
@@
111122222222
Rank Number of singularities
Generic rep. 5 1
1st nongeneric rep. 2 4
2st nongeneric rep. 3 2
4.1. Painlevé I
The standard rank 2 Painlevé I Lax representation has a single irregular singularity, at infinity, given by a pole of order 4, but such that the leading term is nilpotent, which leads to a Turritin-Levelt normal form having a single twisted Stokes circle of slope $5/2$.
The corresponding fission forest (having a single tree) is
5/2525/25 / 222223/2323/23 / $21/$ 2
It admits representatives given by irregular classes of the form
$\bm{\breve{\Theta}}^{I}_{std}=\langle\alpha z^{5/2}\rangle_{\infty},$
for any $\alpha\neq 0$.
Let us study the classes of representations corresponding to the orbit of $\bm{\breve{\Theta}}_{std}^{I}$. The generic class of representations has rank 3, and admits representatives of the form
$\bm{\breve{\Theta}}^{I}_{gen}=\langle\lambda z^{2}+\beta z^{5/3}\rangle_{\infty},$
with $\beta\neq 0$.
The generic tree is
22225/3535/35 / $34/$ 311112/3232/32 / $31/$ 3
There is a single Fourier sphere coefficient here so $N=N_{1}=N_{1}^{+}$, and the standard representation corresponds to the only nongeneric class of representations. Notice that the generic class of representations corresponds to the (abstract version of the more explicit) alternative Lax representation for Painlevé I found in [35].
The corresponding diagram only has one vertex and one loop, and is drawn below:

### v3

@@ 1. $(1)$
    The generic class of representations has rank
    $r_{\rm gen}=\sum_{i}\Bigg(\sum_{j}n^{+}_{ij}r^{+}_{ij}+\sum_{j}n^{-}_{ij}r^{-}_{ij}\Bigg).$
@@ 2. $(2)$
    The $i$-th nongeneric class of representations has rank
    $r_{i}=\sum_{j}n^{+}_{ij}R^{+}_{ij}+\sum_{l\neq i}\Bigg(\sum_{j}n^{+}_{lj}r^{+}_{lj}+\sum_{j}n^{-}_{lj}r^{-}_{lj}\Bigg).$
@@
We can also express the number of singularities at finite distance of the classes of nearby representations.
@@
Let $(E,\nabla)$ of generic form, and $(E^{\prime},\nabla^{\prime})\in\operatorname{SL}_{2}(\mathbb{C})\cdot(E,\nabla)$. If $(E^{\prime},\nabla^{\prime})$ is in the $i$-th nongeneric class of representations, its number of singularities at finite distance is equal to the number of vertices of height 1 in the subtree of $\mathbf{T}$ with set of leaves $N_{i}^{-}$.
@@
By definition of fission trees, keeping previous notations, this number of vertices is equal to the number of distinct coefficients $b_{ij}$ among the Stokes circles of $N_{i}^{-}$. The result is then a direct consequence of item (2) of Proposition 2.16: the position of the singularity above which $F\cdot\bigl\langle q^{-}_{i,j}\bigr\rangle$ lies is determined by the coefficient $b_{ij}$. ∎
In particular, if $N_{i}^{-}=\varnothing$, then the $i$-th nongeneric class of representations only has a singularity at infinity. In general, the class of nearby representations all have different ranks and number of singularities.
Finally, we can also read from the short fission tree when two nongeneric representations have the same fission forest.
@@
Keeping previous notations, for $i,j\in\{1,\dots k\}$, with $i\neq j$, the fission forests $\mathbf{F}_{i}$ and $\mathbf{F}_{j}$ are isomorphic if and only if the $i$-th and $j$-th principal subtrees of the short fission tree $\mathbf{T}$ are isomorphic.
@@
Since isomorphism classes of fission trees characterise admissible deformations, it is equivalent to show that the $i$-th and the $j$-th nongeneric forests are isomorphic if and only if $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are admissible deformations of each other. Let us first assume that $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are admissible deformations of each other. We have to be careful that, in general, if two irregular class $\Theta$ and $\Theta^{\prime}$ at infinity are admissible deformations of each other, this does not imply that their Fourier transforms are also admissible deformations of each other. Indeed, if $\bigl\langle q\bigr\rangle$ is a Stokes circle for $\Theta$, and $\bigl\langle q^{\prime}\bigr\rangle$ is a Stokes circle for $\Theta^{\prime}$ corresponding to $\bigl\langle q\bigr\rangle$ under some admissible deformation, their slopes may differ, and in turn $F\cdot\bigl\langle q\bigr\rangle$ and $F\cdot\bigl\langle q^{\prime}\bigr\rangle$ may not be admissible deformations of each other. However here, because all Stokes circles of $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are of slope <2<2 this situation cannot happen, and we have that $F\cdot\Theta_{i}^{<2}$ and $F\cdot\Theta_{j}^{<2}$ are admissible deformations of each other. In turn, the irregular classes $\Theta_{i}^{<2}+\sum_{l\neq i}\Theta_{l}$ and $\Theta_{j}^{<2}+\sum_{l\neq j}\Theta_{l}$ are admissible deformations of each other, so the $i$-th and $j$-th representations coincide. Conversely, if the $i$-th and $j$-th classes of representations coincide, then in particular $\sum_{l\neq i}\Theta_{l}$ and $\sum_{l\neq j}\Theta_{l}$ are admissible deformations of each other, which implies that $\Theta_{i}$ and $\Theta_{j}$ are admissible deformations of each other, and the conclusion follows. ∎
@@
The classes of nearby representations arise from the action of $\operatorname{SL}_{2}(\mathbb{C})$ on connections on the affine line. However, as already mentioned in the introduction, there is a larger class of natural operations on connections: indeed one can consider Möbius transformations, which have the effect of changing the point at infinity in the identification $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$, as well as the operation consisting in tensoring a connection with an arbitrary rank one connection on a Zariski open subset of $\mathbb{P}^{1}$. This larger class of operations on connections on $\mathbb{P}^{1}$ plays an essential role in the extension due to Deligne–Arinkin [3, 19] of the Katz algorithm for rigid local systems to the case of irregular connections, which provides a way to reduce any irreducible rigid connection on $\mathbb{P}^{1}$ to the trivial rank one connection by repeated application of one of these operations. Since the symplectic $\operatorname{SL}_{2}(\mathbb{C})$ transformations do not commute with these other operations in general, the latter do not preserve the diagram or the classes of nearby representations.
4 Basic representations for Painlevé moduli spaces
@@
In this section, as a set of examples illustrating our general discussion, we describe the basic representations of all Painlevé moduli spaces obtained from the standard rank 2 Lax representation of each Painlevé equation. Recall that, although they were not first discovered in this way historically, all Painlevé equations can be obtained from isomonodromic deformations. Each Painlevé equation arises from isomonodromic deformations of connections with a specific type of singularity data, i.e., different Painlevé equations correspond to different nonabelian Hodge spaces. In the standard way to derive the Painlevé equations from isomonodromy [35], all Lax representations are of rank 2, and in all cases except for Painlevé VI, these Lax representations have at least one irregular singularity.
@@
For each case, starting from this standard Lax representation (in its abstract form, i.e., consisting of the datum of a wild Riemann surface), we will draw the corresponding short fission tree of minimal rank and describe explicitly the classes of nearby representations. While the cases of Painlevé VI, V, IV all fall in the simply-laced case and were discussed by Boalch in [8], it is still worth discussing them briefly from the point of view of short fission trees. On the other hand, describing all basic representations for Painlevé III, II, I as well as for the moduli spaces corresponding to the so-called degenerate versions of Painlevé III really requires using our general framework. Doing this, we recover among the classes of nearby representations many known alternative Lax representations for various Painlevé equations, e.g., from [26, 36, 37] and the other representations can be viewed as new abstract Lax representations.
Remark 4.1.
@@
Notice that there is the following subtlety: since taking the Fourier transform does not always commute with taking admissible deformations, different connections having the same fission forest in general do not have the same classes of nearby representations, nor the same diagram. In the cases where this happens, among all possible wild Riemann surfaces with boundary data having the fission forest of the standard Painlevé Lax representation, we choose one such that its generic representation has minimal rank. Making this choice is necessary to allow us to recover several known alternative Lax representations for Painlevé equations and to recover the diagrams for Painlevé equations obtained in [16], all equal or related to affine Dynkin diagrams. Other choices would lead to different diagrams (typically with more vertices), and for each choice all nearby representations could also be explicitly determined, yielding even more abstract Lax representations, typically of higher rank.
To draw a fission tree $(\mathcal{T},\mathbb{V},\mathbb{A},\mathbb{L},h,n)$, we follow the conventions of [13]. The heights of vertices of the trees are indicated to the left of the figures. The different types of (nonempty) vertices are indicated by different colours: the mandatory vertices, i.e., the elements of $\mathbb{L}$, are drawn in black. The inconsequential vertices, i.e., the elements of $\mathbb{A}\smallsetminus\mathbb{L}$, are drawn in white. Concretely, a black vertex means that in an irregular class with fission tree $\mathcal{T}$ the corresponding coefficient is not allowed to vanish (hence the name mandatory), while a white vertex means that the coefficient is allowed to vanish. Unless otherwise specified, the multiplicity of the leaves in a fission tree is equal to 1.
In [13], fission trees were first defined with an infinite trunk, then suitable truncations of the tree were introduced (see Section 3.5 there) so as to obtain rooted trees with a finite number of vertices: if $\Theta$ is an irregular class, the root of the truncated tree is defined as the (unique) point of the trunk with height $\lfloor\operatorname{Katz}(\Theta)+1\rfloor$.
Here, we will also draw truncated fission trees, and our convention for the height of the root is consistent with loc. cit., as follows. If $\mathcal{F}$ is the fission datum of some irregular class, we define its minimal Poincaré–Katz rank as
$\operatorname{Katz}_{\min}(\mathcal{F}):=\min\{\operatorname{Katz}(\Theta)\mid\Theta\text{ irregular class with }\mathcal{F}(\Theta)=\mathcal{F}\}.$
In turn, the height of the root of the fission tree $\mathbf{T}$ corresponding to $\mathcal{F}$ is taken to be
$\eta:=\lfloor\operatorname{Katz}_{\min}(\mathcal{F})+1\rfloor.$
@@
To draw the diagrams, we follow the conventions of [20]: unless otherwise specified the multiplicity of each vertex is 1, negative edges or loops are drawn in dashed lines, and for multiple edges/loops the multiplicity is indicated at their middle. Furthermore, here we use colours to draw the vertices, each colour corresponding to a subset in the partition defined by the short fission tree.
4.1 Summary of basic Painlevé representations
@@
The short fission trees as well as the rank, number of singularities, and minimal Poincaré–Katz ranks of each class of basic representations for each Painlevé moduli space are listed below. In the next section, we will explain in more detail how to obtain these classes of representations starting from the standard Lax representations.
@@
5/35/34/34/3112/32/31/31/3
Rank Number of singularities min. Katz ranks
Generic rep. 3 1 5/3
Nongeneric rep. 2 1 5/2
@@
11223/23/21/21/2
Rank Number of singularities min. Katz ranks
Generic rep. 3 1 2
1st nongeneric rep. 2 1 3
2nd nongeneric rep. 2 2 3/2, 0
@@
11221/21/2
Rank Number of singularities min. Katz ranks
Generic rep. 4 1 2
1st nongeneric rep. 2 3 1/2, 0, 0
2nd nongeneric rep. 2 2 1, 1
@@
11221/21/2
Rank Number of singularities min. Katz ranks
Generic rep. 4 1 2
Nongeneric reps. 2 2 1, 1/2
@@
11221/21/21/31/3
Rank Number of singularities min. Katz ranks
Generic rep. 5 1 2
1st nongeneric rep. 2 2 1/2, 1/2
2nd nongeneric rep. 3 2 1, 1/3
@@
2211
Rank Number of singularities min. Katz ranks
Generic rep. 3 1 2
Nongeneric reps. 2 2 2, 0
@@
1122
Rank Number of singularities min. Katz ranks
Generic rep. 4 1 2
Nongeneric reps. 2 3 1, 0, 0
@@
112222
Rank Number of singularities min. Katz ranks
Generic rep. 5 1 2
1st nongeneric rep. 3 2 1, 0
2nd nongeneric rep. 2 4 0, 0, 0, 0
4.2 Painlevé I
4.2.1 Standard representation
The standard rank 2 Painlevé I Lax representation has a single irregular singularity given by a pole of order 4, but such that the leading term is nilpotent, which leads to a Turritin–Levelt normal form having a single twisted Stokes circle of slope $5/2$.
This corresponds to the following fission forest $\mathbf{F}^{\rm I}_{\rm std}$ (having a single tree):
5/25/2223/23/2111/21/2
The global irregular classes with fission forest $\mathbf{F}^{\rm I}_{\rm std}$ and minimal Poincaré–Katz rank (equal to $5/2$) are exactly those of the form
$\bm{{\Theta}}^{\rm I}_{\rm std}=\bigl\langle\alpha_{5/2}z_{a}^{-5/2}+\alpha_{2}z_{a}^{-2}+\alpha_{3/2}z_{a}^{-3/2}+\alpha_{1}z_{a}^{-1}+\alpha_{1/2}z_{a}^{-1/2}\bigr\rangle_{a},$
with $a\in\mathbb{P}^{1}$, $\alpha_{5/2},\alpha_{2},\alpha_{3/2},\alpha_{1},\alpha_{1/2}\in\mathbb{C}$, and $\alpha_{5/2}\neq 0$.
Given an irregular connection on $\mathbb{P}^{1}$ with such an irregular class, the rank of its generic class of nearby representations is minimal when $a=\infty$ (to which we can always reduce up to applying a Möbius transformation).
Let us thus consider a (modified) irregular class of the form
$\bm{\breve{\Theta}}^{\rm I}_{\rm std}=\bigl\langle\alpha_{5/2}z^{5/2}+\alpha_{2}z^{2}+\alpha_{3/2}z^{3/2}+\alpha_{1}z+\alpha_{1/2}z^{1/2}\bigr\rangle_{\infty},$
with $\alpha_{5/2}\neq 0$, and determine the classes of its nearby representations.
4.2.2 Generic representation
The short fission tree $\mathbf{T}^{\rm I}$ of $\bm{\breve{\Theta}}^{\rm I}_{\rm std}$ is the following:
5/35/34/34/3112/32/31/31/3
An irregular class of generic form has fission tree $\mathbf{T}^{I}$ if and only if it is of the form
$\bm{\breve{\Theta}}^{\rm I}_{\rm gen}=\bigl\langle\lambda^{\prime}z^{2}+\alpha^{\prime}_{5/3}+\alpha^{\prime}_{4/3}z^{4/3}+\alpha^{\prime}_{1}z+\alpha^{\prime}_{2/3}z^{1/3}+\alpha^{\prime}_{1/3}z^{1/3}\bigr\rangle_{\infty},$
with $\alpha^{\prime}_{5/3}\neq 0$.
Remark 4.2.
Notice that the coefficient $\lambda^{\prime}$ of exponent 2 may vanish, so the minimal Poincaré–Katz rank for such an irregular class is $5/3$, hence following our conventions for truncations of fission trees, we draw the root of the tree at height $\big\lfloor\frac{5}{3}+1\big\rfloor=2$.
The generic class of representations thus has rank 3, a single Fourier sphere coefficient, so $N=N_{1}=N_{1}^{+}$, and the standard representation corresponds to the sole nongeneric class of representations.
Notice that the generic class of representations corresponds to the (abstract version of the more explicit) alternative Lax representation for Painlevé I found in [37].
4.2.3 Diagram
The corresponding diagram $\Gamma^{I}$ has one vertex with one loop [20, Section 7], as drawn below